Trace the curve y2=(x-a)(x-b)(x-c) with a, b, and c are all positive.  


1
Expert's answer
2021-12-28T15:43:33-0500

We consider the following cases:

Case I : "a < b < c"

(1) It is symmetrical about the "x" -axis.

(2) It meets the "x" -axis in"(a, 0), (b,0)" and "(c, 0)."

(3) When "x < a," "y^2" is negative,

when "a<x<b, y^2>0"

when "b<x<c, y^2" is negative,

when "x>c, y^2>0."

Hence, there is no curve to the left-of "x =a" and also between "x =b" and "x=c."

(4) If "x > c" and increases then "y^2" also increases.




Case II : "a = b < c"


"y^2=(x-a)^2(x-c)"


(1) It is symmetrical about the "x" -axis.

(2) It meets the "x" -axis in"(a, 0)" and "(c, 0)."

(3) When "x < a," "y^2" is negative,

when "a<x<c, y^2" is negative.

"(a, 0)" is isolated point.

(4) If "x > c" and increases then "y^2" also increases.



Case III : "a < b = c"


"y^2=(x-a)(x-b)^2"

(1) It is symmetrical about the "x" -axis.

(2) It meets the "x" -axis in"(a, 0)" and "(b, 0)."

(3) When "x < a," "y^2" is negative,

(4) If "x > b" and increases then "y^2" also increases.




Case IV : "a = b = c"


"y^2=(x-a)^3"

(1) It is symmetrical about the "x" -axis.

(2) It meets the "x" -axis in"(a, 0)."

(3) When "x < a," "y^2" is negative,

(4) If "x >a" and increases then "y^2" also increases.